JACKKNIFE VARIANCE ESTIMATORS OF THE LOCATION ESTIMATOR IN THE ONE-WAY RANDOM-EFFECTS MODEL

YOUNGJO LEE

Department of Statistics, Hallym University, Chunchon, Korea

(Received October 11, 1989; revised July 6, 1990)

Abstract. In a one-way random-effects model, we frequently estimate the variance components by the analysis-of-variance method and then, assuming the estimated values are true values of the variance components, we estimate the population mean. The conventional variance estimator for the estimate of the mean has a bias. This bias can become severe in contaminated data. We can reduce the bias by using the delta method. However, it still suffers from a large bias. We develop a jackknife variance estimator which is robust with respect to data contamination.

Key words and phrases: Jackknife, one-way random-effect models, robustness, Monte Carlo study.